this is the word problem. The profit function for the Recklus Hang gliding Service is P(x) = -0.2x
2
+ fx - m,
where f represents the set up fee for a customer’s daily excursion and m
represents the monthly hanger rental. Also, P represents the monthly profit in
dollars of the small business where x is the number of flight excursions facilitated
in that month.
The question i have is How many flight excursions must be sold in order to maximize the profit? Show
your work algebraically. Trial and error is not an appropriate method of solution –
use methods taught in class

Answer by stanbon(75887)

The profit function for the Recklus Hang gliding Service is
P(x) = -0.2x^2 + fx - m, where f represents the set up fee
for a customer’s daily excursion and m represents the monthly
hanger rental.
----------------------
Also, P represents the monthly profit in dollars of the small
business where x is the number of flight excursions facilitated
in that month.
The question i have is How many flight excursions must be sold in order to maximize the profit?
--------------------
Show your work algebraically.
Trial and error is not an appropriate method of solution –
---
Maximum profit occurs when x = -b/2a = -f/(2*-0.2) = f/0.4
--------------
Cheers,
Stan H.